the average price of an acre of U.S. farmland was $1210 in 2002. in 2008, the price of an acre rose to $2350.
a. write two ordered pairs of the form
b. find the slope of the line through two points
c.write a sentence explaining the meaning of the slope as a rate of change

Answer by fractalier(6550)

a) (2002, 1210) and (2008, 2350) (Usually use x for time)
b) slope = (2350-1210)/(2008-2002) = 1140/6 = $190/yr
c) The price of an acre of land had an average increase of $190 per year between the years 2002 and 2008.